Character’s Name: Abel Marko  ^{Player’s Name: chad brown}^{Race:}  ^{Human}  ^{Class:}  ^{Cleric^{Priest}}  ^{AL:} ^{LG}  ^{Lvl:} ^{17}  ^{Sex:} ^{M}^{Weight:} ^{245 lb.}  ^{Age:}  ^{25 yrs}  ^{Height:} ^{6’4” in.}^{App.:}  big, BIG, BIG, muscular^{Subrace:}  ^{Religion:} ^{God}  ^{Eyes:} ^{Green}  ^{Hair:} ^{bald}^{Children:}^{none}  ^{Siblings:} ^{brother, Cain}  ^{Homet.:} ^{Delic}^{Backg.:} ^{Member of Justice}

STR: 16 DEX: 16 CON: 15 INT: 14 WIS: 18 CHR: 15

{Hit. \atop Prob.}  {Dmg. \atop Adj..}  {Wgt. \atop All.}  Max Press  {Op. \atop Drs.}  {B.B/ \atop L.G.}
0  1  70  195  9  10%
{Rctn. \atop Adj.}  Missile Att Adj.  {Def. \atop Adj.}
1  1  -2
{HP. \atop Adj.}  {Sys. \atop Shk.}  {Res. \atop Sur.}  {Pois. \atop Sav.}  Reg.
1  90%  94%  0  Nil
No.of. Lang.  Spell Lvl.  {Lrn. \atop Sp.}  {Sp/ \atop Lvl.}  {Sp. \atop Imm.}
4  7th  60%  9  0. Ill
{Mg.D. \atop Adj.}  {Bon. \atop Sp.}  {Sp. \atop Fail.}  {Sp. \atop Imm.}
4  Yes  0%  No
Max. Hench.  {Loy. \atop Base}  {Rctn. \atop Adj.}
7  3  3

Movement^{Base:}^{Lgh:}^{Mod:}^{Hvy:}^{Sev:}^{Encumbrance:} ^{12}^{9}^{(71 - 100)} ^{6}^{(101 - 130)} ^{3}^{(131 - 160)} ^{1}^{(161 - 195)}

Mod. ^{Sav.Throw} ^{Base} 10 4 8 9 7^{Breath}^{Weapon}^{Paralyze}^{Poison}^{Rod,Staff}^{or Wand}^{Spells}^{Petrify}^{Polym.}

AC 9 ^{Modifiers}  ^{Armor}  ^{Weight}^{Surprised}^{Shieldless}^{Rear}  ^{5}^{5}^{5} ^{Robes}  ^{1 lb.}
HP 82^{Base: 106}  ^{Wounds}

Weapon  ^{#Att} ^{AttAdj} ^{DAdj} ^{THAC0} ^{DmgSM} ^{DmgL}  ^{Rg S}  ^{Rg M}  ^{Rg L}  ^{Wght}  ^{Type}  ^{Size}  ^{Speed} Dagger  2  1  10 1d4 1d3  1 lb. P  S  7

Proficiencies

Languages:

common elvish dwarvish orc

Weapon:

ambidexterity two-weapon style Club long sword long bow lance Glaive dirk

Non-Weapon:

Riding, Land-Based Swimming Animal Training Religion Spellcraft Healing (W/+3) (S/0) (W/0) (W/0) (I/-2) (I/-2) (W/-2) (I/+1) (NA/NA)

( - )

( - )

( - )

Special Attacks Martial Arts  ^{Ammunition}

Abilities

Herbalism Reading/Writing Blind fighting

( - )

( - )

Gear^{Location} ^{Item}  ^{Weight} Backpack  2 lb. Picks  1 lb. 9 Candles  1 lb. Clear Crystal  1 lb. Robe  1 lb. Wineskin  1 lb. Rope (50ft)  1 lb.  17 Sheet of paper  1 lb.^{Location} ^{Item}  ^{Weight}

Water Wine Rations

Feed

XP^{Bonus: 20%}

Gems^{8 rubies}  ^{Coins} 220^{cp} 22^{sp}  2210 ^{ep} 121 ^{gp}  1330 ^{pp}  ^{other valuabels} 5 very valuable Art Objects

Misc. Information (Magical Items, Command Words, Small Maps etc.)

slide rule

Henchmen/Animal Companions Name  AC  Race/Class  HD/Lvl #AT THAC0  Dmg/Effects  Skills/Abilities  HP Background: Background: Background: Background: Background: Background: Background: Background: